A projectile is thrown with a velocity of 10Root2 m/s at an angle of 45 degrees with horizontal. The interval between the moments when speed is root125 m/s 

Grade:9

1 Answers

Arun
3 years ago
Need to find : The time interval between the moments when the speeds are √125 m/s
 
projectile is thrown with a velocity of 10√2 m/s
 
angle of 45°
 
Horizontal Velocity = 10√2Cos45  = 10 m/s
 
Vertical Velocity = 10√2Sin45  = 10 m/s
 
Horizontal Velocity will remain constant = 10
 
Let say at time t the Velocity = √125 m/s
 
=> Vertical Velocity² + Horizontal Velocity² =( √125 )²
 
=> Vertical Velocity²  + 10² = 125
 
=> Vertical Velocity² = 25
 
=> Vertical Velocity = ±5
 
Applying V = U + at
 
5 = 10 + (-10)t
 
=> t = 0.5 sec
 
-5 = 10  + (-10)t
 
=> t = 1.5 sec
 
√125 m/s velocity would be at t =0.5 sec & 1.5 sec
 
Time interval between = 1.5  - 0.5 = 1 sec